# Perspective: Interactions mediated by atoms, photons, electrons, and excitons

### v1

@@ II Interactions between quasiparticles
In this section, we will set up the general framework for analysing the interactions between quasiparticles as first put forward by Landau, and which is a pinnacle in theoretical physics allowing a relatively simple yet powerful description of otherwise complicated interaction quantum many-body systems. The premise of Landau’s theory is that the energy of a many-body system can be expanded as a power series in the occupation numbers of its fundamental excitations Baym and Pethick (2008), which behave as particle-like entities, so-called quasiparticles, with renormalised properties and finite life-times. Keeping terms up to second order in the occupation numbers, the energy of the system is written as
@@ III.1 Fermion mediated interaction
We will now calculate the interaction between two Fermi polarons assuming the interaction $g$ is weak so that perturbation theory applies. An eigenstate for no interactions ($g=0$) is $|\psi_{0}\rangle=\prod_{\mathbf{p}}\hat{a}^{\dagger}_{\mathbf{p}}\prod_{|\mathbf{k}|<k_{F}}\hat{b}^{\dagger}_{\mathbf{k}}|0\rangle$ where $|0\rangle$ is the vacuum state and $k_{F}$ is the Fermi momentum. The energy shift to first order is $E_{1}=\langle\psi_{0}|\hat{H}_{\text{int}}|\psi_{0}\rangle=N_{a}gn_{F}$, which is simply the mean-field energy coming from the density $n_{F}$ of the fermions. The second order energy shift is $E_{2}=\sum_{n\neq 0}\left|\langle\psi_{0}|\hat{H}_{\text{int}}|\psi_{n}\rangle\right|^{2}/(E_{0}-E_{n})$. The relevant unperturbed eigenstates are $|\psi_{n}\rangle=[n_{\mathbf{p}a}(n_{\mathbf{p}-\mathbf{q}a}+1)]^{-1/2}\hat{a}^{\dagger}_{\mathbf{p}-\mathbf{q}}\hat{a}_{\mathbf{p}}\hat{b}^{\dagger}_{\mathbf{k}+\mathbf{q}}\hat{b}_{\mathbf{k}}|\psi_{0}\rangle$ where $n_{\mathbf{p}a}$ is the occupation function of the impurities. This straightforwardly gives
@@
where $n_{\mathbf{k}b}=1/\{\exp[(\epsilon_{\mathbf{k}b}-\mu_{b})/T]+1\}^{-1}$ is the occupation function for the fermions with chemical potential $\mu_{b}$. The upper/lower sign in the second bracket in Eq. (6) is for bosonic/fermionic impurities and originates from Bose stimulation/Fermi blocking of the scattering processes.
@@ III.2 Boson mediated interaction
Physically, the $1/g_{b}$ dependence in Eq. (13) reflects that the interaction increases with the compressibility of the BEC. Like for the fermionic case discussed in Sec. III.1, this derivation explicitly demonstrates the key role of quantum statistics making the quasiparticle interaction attractive/repulsive for bosonic/fermionic quasiparticles.
@@ IV.1 Green’s function formalism
Green’s functions provide a compact way to analyse interacting many-particle systems Fetter and Walecka (1971); Abrikosov et al. (2012). Consider again the $N_{a}$ impurity particles of mass $m_{a}$ in a medium of majority particles. The Green’s function of an impurity with momentum $\mathbf{p}$ is given by $G_{a}(\mathbf{p},\omega)$ and it is governed by the Dyson’s equation $G_{a}^{-1}(\mathbf{p},\omega)=G^{-1}_{a0}(\mathbf{p},\omega)-\Sigma_{a}(\mathbf{p},\omega)$ where $G^{-1}_{a0}(\mathbf{p},\omega)=\omega-\epsilon_{\mathbf{p}a}$ is the noninteracting Green’s function and $\Sigma_{a}(\mathbf{p},\omega)$ the self-energy of the impurity atoms. The energy of the quasiparticle state of an impurity with momentum $\bf p$ is determined by the self-consistent equation Fetter and Walecka (1971); Abrikosov et al. (2012)
@@
$\Sigma_{2a}({\bf p},i\omega_{p})=\pm g^{2}\int\!\frac{d^{3}p^{\prime}}{(2\pi)^{3}}\chi({\bf p}-{\bf p}^{\prime},i\omega_{p})n_{{\bf p}^{\prime}a},$ (17)
@@ IV.2 Path integral approach
From this we see that the BEC mediates a frequency dependent interaction $g^{2}\chi(\bf p,\omega)$ (see Eq. (12)) between the impurities with the compressibility of the BEC given by Eq. (12). Calculating the second order Fock diagram with this interaction then gives Eq. (17) from which Eq. (7) for the quasiparticle interaction again is recovered.
@@ VII.1 Mediated interactions in 3D
We see that the energy shift is negative for a bosonic impurity (left) corresponding to an attractive polaron-polaron interaction, whereas it is positive for a fermionic impurity (right) corresponding to a repulsive interaction consistent with the results in Sec. III.2. Even though such energy shifts should be detectable, the quasiparticle interaction between Bose polarons remains to be observed, possibly due to the generally broad experimentally energy spectra. It has been shown that an alternative way to observe the BEC mediated interaction is via the sizable frequency shift it causes for the out-of-phase dipole mode of two Fermi gases separated by a BEC as compared to their in-phase mode Suchet et al. (2017)
Using the interaction given Eq. (27) in an effective Schrödinger equation for two polarons, it was predicted to be strong enough to support bound states, i.e. bi-polarons Camacho-Guardian et al. (2018). The results for the binding energy of the bi-polaron were furthermore shown to compare well with Monte-Carlo calculations, and to recover exact results for the bound states of the Yukawa interaction given by Eq. (23) for weak interactions. The same approach was also used to predict bound states between two Bose polarons both below and above the two-particle continuum in a 2D square lattice due to an attractive quasiparticle interaction Ding et al. (2023). Within the Fröhlich model, valid for weak interactions, bipolarons were furthermore predicted using an all-coupling treatment in Ref. Casteels et al. (2013). The problem of a gas of diluite polarons has been studied recently with perturbative a field-theory and a quantum–Monte Carlo (QMC) method in Ref. Ardila (2022).
@@
Induced interactions are not limited to the case where the impurity-boson interaction is short-ranged. For instance, progress creating hybrid ion-atom systems Grier et al. (2009); Härter et al. (2012); Kleinbach et al. (2018); Tomza et al. (2019); Dieterle et al. (2021) has stimulated much attention to charged polarons Casteels et al. (2011); Astrakharchik et al. (2021); Christensen et al. (2021, 2022); Pessoa et al. (2024) and their interactions. The interaction between two static ions immersed in a neutral atomic BEC or Fermi gas was considered in Ref. Ding et al. (2022). For weak interactions, it was shown that the mediated between two ions mediated by a BEC exhibits a Yukawa form for intermediate distances and a $1/r^{4}$ power-law behaviour for large distances reminiscent of the ion-atom interactions. For strong atom-ion interactions, a calculation similar to that leading to Eq. 27 shows that for large distances the mediated interaction is proportional to $\mathcal{T}(0,0;0)/r^{4}$ where $\mathcal{T}(0,0;0)$ is the zero energy/momentum atom-ion scattering matrix. For two static ions interacting weakly with a Fermi gas, the mediated interaction is given by a power-law for large density and by a RKKY form for low density. Using Monte Carlo simulations, the interaction between two static impurities in a weakly interacting BEC was studied Astrakharchik et al. (2023). The ions were found to strongly perturb the BEC in their surroundings and different regimes related to the formation of bi-polarons were identified.
@@ VII.2 Mediated interactions in 1D
This can be derived from a Fourier transform of the 1D version of Eq. 12 and corresponds to the exchange of a singly phonon between the impurities as illustrated in Fig. 7 (b). For large distances $r\gg\xi$ where the two dressing clouds do not overlap, the interaction between the impurities is due to quantum fluctuations in close analogy with the Casimir effect as already discussed in Sec. VI. This gives rise to the interaction Schecter and Kamenev (2014); Pavlov et al. (2019); Reichert et al. (2019, 2019)
@@
exhibiting the same $1/r^{3}$ scaling as the Casimir force. It is caused by the exchange of two phonons as illustrated in Fig. 7 (c). Further studies concerning polaron interactions and bipolarons in 1D emphasised the deformation of the superfluid by the impurities Will et al. (2021) and the presence of lattice 1D bipolarons Isaule et al. (2024).
Figure 7: Figure from Ref. Schecter and Kamenev (2014). (a) The interaction between two static impurities mediated by a 1D Bose gas. (b) Single-phonon exchange responsible for the exponential decay of the interaction at short distances. (c) Two-phonon exchange giving rise to a long-range Casimir-like interaction between the impurities. The solid lines correspond to the impurity propagator while the wavy line illustrate the phonon modes.
@@ VII.3.1 General Theoretical framework
Figure 9: Diagrammatic representation of the photon mediated interaction between ground state atoms to the lowest order in $\Delta_{a}$, which involves the exchange of a single photon. The solid yellow/red lines depict the atomic ground/excited state Green’s functions, the crossed dot with the black wiggly line corresponds to the classical pump field, the white dot corresponds to the coupling to the cavity photons, and the wavy blue line is the photon propagator.
@@
First, our main assumption is that the pump and cavity frequencies are far detuned from the excited state such that $\Delta_{a}$ is large. This is the so-called dispersive regime where the population of the excited state is small and one can write their Green’s function as $G_{e}(k)\simeq 1/\Delta_{a}$. The leading diagram in $1/\Delta_{a}$ for the photon-mediated interaction between the ground state atoms can then be drawn as in Fig. 9(a), which intuitively can be interpreted from left to right as: a ground state atom (orange line) is excited to the excited state (red line) by the classical pump field $\Omega$ (red cross). The excited state then decays to the ground state again by emitting a photon into cavity mode (blue dashed line). The cavity photon is subsequently absorbed by a second ground state atom, which is then excited. This atom finally decays back into the ground state atom via stimulated emission into the classical field. This completes the mediated interaction, as two ground state atoms effectively exchange a photon. As illustrated in Fig. 9(b), the processes highlighted by the box can be summarized into an effective interaction $\eta({\mathbf{r}})=\Omega({\mathbf{r}})\mathcal{G}({\mathbf{r}})/\Delta_{a}$ between a ground state atom and a photon. Thus, as for the other mediated interactions discussed in this perspective, the photon mediated interaction between the ground state atoms is quadratic in $\eta$. Finally, assuming that the dynamics of the cavity photon can be ignored when $\Delta_{c}$ is large so that its Green’s function is $G_{c}(k)\simeq 1/\Delta_{a}$ and taking into account the two processes in Fig. 9(c) gives
@@
for the photon mediated interaction between the ground state atoms. Note that this interaction is infinite range reflecting that the cavity mode is assumed to cover the whole system. The photon-mediated interactions can be derived alternatively using a path-integral formalism as in Sec. IV.2. This approach has been employed for the study of supersolid phases of bosonic atoms inside two crossed optical cavities Lang et al. (2017); Piazza et al. (2013).
@@
$\left.+g_{eg}{\hat{\psi}}_{g}^{\dagger}(\mathbf{r}){\hat{\psi}}_{g}(\mathbf{r})\right]{\hat{\psi}}_{g}(\mathbf{r})+\hbar\left[\Omega(\mathbf{r})+\mathcal{G}(\mathbf{r})\hat{a}^{\dagger}\right]{\hat{\psi}}_{e}(\mathbf{r})$ (34)
@@
Here, we have made the identifications, $\partial{\hat{\psi}}_{g}(\mathbf{r})=\partial{\hat{\psi}}(\mathbf{r})$, $V({\mathbf{r}})=\Omega^{2}({\mathbf{r}})/\Delta_{a}$, $U({\mathbf{r}})=\mathcal{G}({\mathbf{r}})/\Delta_{a}$, $\hat{n}(\mathbf{r})={\hat{\psi}}^{\dagger}_{g}(\mathbf{r}){\hat{\psi}}_{g}(\mathbf{r})$, and $g_{gg}=g_{0}$. The above equations establish the coupled dynamics between matter and light.
@@
We now consider the case in which the cavity field dynamics evolves much faster than the atom center of mass motion, which is equivalent to the assumption that $\Delta_{c}$ is large used in the diagrammatic derivation. Then, the cavity field follows adiabatically the movement of the atoms, as in the Born-Oppenheimer approximation scheme of electrons in molecules and using the steady-state solution to the photonic field operator $\partial\hat{a}/\partial t=0$ allows one to derive a ground state atom only description of the system with a photon mediated interaction given by Eq. (32).

### v3

@@ II Interactions between quasiparticles
In this section, we will set up the general framework for analysing the interactions between quasiparticles as first put forward by Landau, and which is a pinnacle in theoretical physics allowing a relatively simple yet powerful description of otherwise complicated interaction quantum many-body systems. The premise of Landau’s theory is that the energy of a many-body system can be expanded as a power series in the occupation numbers of its fundamental excitations Baym and Pethick (2008), which behave as either bosonic or fermionic particle-like entities, so-called quasiparticles, with renormalised properties, finite life-times. Keeping terms up to second order in the occupation numbers, the energy of the system is written as
@@ III.1 Fermion mediated interaction
Before we delve into the calculations, we note that while Eq. 5 relates to the interaction between a bare impurity and a fermion from the Fermi sea, Landau’s quasiparticle relates to an effective interaction between quasiparticles, which we now calculate. We begin with the interaction between two Fermi polarons assuming the interaction $g$ is weak so that perturbation theory applies. An eigenstate for no interactions ($g=0$) is $|\psi_{0}\rangle=\prod_{\mathbf{p}}\hat{a}^{\dagger}_{\mathbf{p}}\prod_{|\mathbf{k}|<k_{F}}\hat{b}^{\dagger}_{\mathbf{k}}|0\rangle$ where $|0\rangle$ is the vacuum state and $k_{F}$ is the Fermi momentum. The energy shift to first order is $E_{1}=\langle\psi_{0}|\hat{H}_{\text{int}}|\psi_{0}\rangle=N_{a}gn_{F}$, which is simply the mean-field energy coming from the density $n_{F}$ of the fermions. The second order energy shift is $E_{2}=\sum_{n\neq 0}\left|\langle\psi_{0}|\hat{H}_{\text{int}}|\psi_{n}\rangle\right|^{2}/(E_{0}-E_{n})$. The relevant unperturbed eigenstates are $|\psi_{n}\rangle=[n_{\mathbf{p}a}(n_{\mathbf{p}-\mathbf{q}a}+1)]^{-1/2}\hat{a}^{\dagger}_{\mathbf{p}-\mathbf{q}}\hat{a}_{\mathbf{p}}\hat{b}^{\dagger}_{\mathbf{k}+\mathbf{q}}\hat{b}_{\mathbf{k}}|\psi_{0}\rangle$ where $n_{\mathbf{p}a}$ is the occupation function of the impurities. This straightforwardly gives
@@
where $n_{\mathbf{k}b}=1/\{\exp[(\epsilon_{\mathbf{k}b}-\mu_{b})/T]+1\}^{-1}$ is the occupation function for the fermions with chemical potential $\mu_{b}$. The upper/lower sign in the second bracket in Eq. (6) is for bosonic/fermionic impurities and originates from Bose stimulation/Fermi blocking of the scattering processes. Mathematically, it arises from the (anti)-commutation rules for (fermions) and bosons.
@@ III.2 Boson mediated interaction
Physically, the $1/g_{b}$ dependence in Eq. (13) reflects that the interaction increases with the compressibility of the BEC. The second-order contribution in Eq. III.2 arising from the quartic term in Eq. 10 is suppressed by the gas factor $n_{0}a^{3}\ll 1$  Levinsen et al. (2017) and gives a small correction to the induced interaction, except for very low momenta where it gives an unphysical divergence for $\lim_{|\mathbf{p}|\rightarrow|\mathbf{p}^{\prime}|}f_{\mathbf{p},\mathbf{p}^{\prime}}.$
Like for the fermionic case discussed in Sec. III.1, this derivation explicitly demonstrates the key role of quantum statistics making the quasiparticle interaction attractive/repulsive for bosonic/fermionic quasiparticles.
@@ IV.1 Green’s function formalism
Green’s functions provide a compact way to analyse interacting many-particle systems Fetter and Walecka (1971); Abrikosov et al. (2012). Consider again the $N_{a}$ impurity particles of mass $m_{a}$ in a medium of majority particles. The retarded Green’s function of an impurity with momentum $\mathbf{p}$ is given by $G_{a}(\mathbf{p},\omega)$ and it is governed by the Dyson’s equation $G_{a}^{-1}(\mathbf{p},\omega)=G^{-1}_{a0}(\mathbf{p},\omega)-\Sigma_{a}(\mathbf{p},\omega)$ where $G^{-1}_{a0}(\mathbf{p},\omega)=\omega-\epsilon_{\mathbf{p}a}$ is the noninteracting Green’s function and $\Sigma_{a}(\mathbf{p},\omega)$ the self-energy of the impurity atoms. The energy of the quasiparticle state of an impurity with momentum $\bf p$ is determined by the self-consistent equation Fetter and Walecka (1971); Abrikosov et al. (2012)
@@
$\Sigma_{2a}({\bf p},i\omega_{p})=\pm g^{2}\int\!\frac{d^{3}p^{\prime}}{(2\pi)^{3}}\chi({\bf p}-{\bf p}^{\prime},\omega)n_{{\bf p}^{\prime}a},$ (17)
@@ IV.2 Path integral approach
From this we see that, after analytic continuation $i\omega_{p}\rightarrow\omega+i0^{+}$, the BEC mediates a frequency dependent interaction $g^{2}\chi(\bf p,\omega)$ (see Eq. (12)) between the impurities with the compressibility of the BEC given by Eq. (12). Calculating the second order Fock diagram with this interaction then gives Eq. (17) from which Eq. (7) for the quasiparticle interaction again is recovered. Here again, we ignore the quartic term in the impurity-boson coupling, which is expected to give small corrections.
@@ VI.2 Mediated interactions in 1D
Interestingly, the theoretical framework used to study Fermi polarons has found direct applications in nuclear physics, such as cluster formation in neutron-rich matter Tajima et al. (2024a); Guo and Tajima (2024); Tajima et al. (2024b).
@@ VII.1 Mediated interactions in 3D
We see that the energy shift is negative for a bosonic impurity (top) corresponding to an attractive polaron-polaron interaction, whereas it is positive for a fermionic impurity (bottom) corresponding to a repulsive interaction consistent with the results in Sec. III.2. Even though such energy shifts should be detectable, the quasiparticle interaction between Bose polarons remains to be observed, possibly due to the generally broad experimentally energy spectra. It has been shown that an alternative way to observe the BEC mediated interaction is via the sizable frequency shift it causes for the out-of-phase dipole mode of two Fermi gases separated by a BEC as compared to their in-phase mode Suchet et al. (2017)
Using the interaction given Eq. (27) in an effective Schrödinger equation for two polarons, it was predicted to be strong enough to support bound states, i.e. bi-polarons Camacho-Guardian et al. (2018). The results for the binding energy of the bi-polaron were furthermore shown to compare well with Monte-Carlo calculations, and to recover exact results for the bound states of the Yukawa interaction given by Eq. (23) for weak interactions. The same approach was also used to predict bound states between two Bose polarons both below and above the two-particle continuum in a 2D square lattice due to an attractive quasiparticle interaction Ding et al. (2023). Within the Fröhlich model, valid for weak interactions, bipolarons were furthermore predicted using an all-coupling treatment in Ref. Casteels et al. (2013). The problem of a gas of diluite polarons has been studied recently with perturbative a field-theory and a quantum–Monte Carlo (QMC) method in Ref. Ardila (2022). Squeezing and entanglement of distinguishable polarons have been studied in the context of two quantum Brownian particles coupled to a bosonic bath Charalambous et al. (2019).
@@
Induced interactions are not limited to the case where the impurity-boson interaction is short-ranged. For instance, progress creating hybrid ion-atom systems Grier et al. (2009); Härter et al. (2012); Kleinbach et al. (2018); Tomza et al. (2019); Dieterle et al. (2021) has stimulated much attention to charged polarons Casteels et al. (2011); Astrakharchik et al. (2021); Christensen et al. (2021, 2022); Pessoa et al. (2024) and their interactions. The interaction between two static ions immersed in a neutral atomic BEC or Fermi gas was considered in Ref. Ding et al. (2022). For the weakly interacting case, it was shown that the interaction between two ions mediated by a BEC, exhibits a Yukawa form for intermediate distances, and a $1/r^{4}$ power-law behaviour for large distances, reminiscent of the ion-atom interactions. This Yukawa interaction for long distances arises also for the mediated interaction between Rydberg impurities Wang et al. (2015). For strong atom-ion interactions, a calculation similar to that leading to Eq. 27 shows that for large distances the mediated interaction is proportional to $\mathcal{T}(0,0;0)/r^{4}$ where $\mathcal{T}(0,0;0)$ is the zero energy/momentum atom-ion scattering matrix. For two static ions interacting weakly with a Fermi gas, the mediated interaction is given by a power-law for large density and by a RKKY form for low density. Using Monte Carlo simulations, the interaction between two static impurities in a weakly interacting BEC was studied Astrakharchik et al. (2023). The ions were found to strongly perturb the BEC in their surroundings and different regimes related to the formation of bi-polarons were identified.
@@ VII.2 Mediated interactions in 1D
This can be derived from a Fourier transform of the 1D version of Eq. 12 and corresponds to the exchange of a single phonon between the impurities as illustrated in Fig. 7 (b). This diagram corresponds to the one described in detail using second-order perturbation theory in Fig. 2(c), where impurities exchange a Bogoliubov mode from the condensate. For large distances $r\gg\xi$ where the two dressing clouds do not overlap, the interaction between the impurities is due to quantum fluctuations in close analogy with the Casimir effect as already discussed in Sec. VI. This gives rise to the interaction Schecter and Kamenev (2014); Pavlov et al. (2019); Reichert et al. (2019, 2019)
@@
exhibiting the same $1/r^{3}$ scaling as the Casimir force. It is caused by the exchange of two phonons as illustrated in Fig. 7 (c). Further studies concerning polaron interactions and bipolarons in 1D emphasised the deformation of the superfluid by the impurities Will et al. (2021); Brauneis et al. (2021), the presence of lattice 1D bipolarons Isaule et al. (2024), and bipolarons maximally entangled appearing in the form of Bell states Anh-Tai et al. (2024).
Figure 7: Figure from Ref. Schecter and Kamenev (2014). (a) The interaction between two static impurities mediated by a 1D Bose gas. (b) Single-phonon exchange responsible for the exponential decay of the interaction at short distances. Indeed, this process corresponds to the one discussed in Fig. 2(c). (c) Two-phonon exchange giving rise to a long-range Casimir-like interaction between the impurities. The solid lines correspond to the impurity propagator while the wavy line illustrate the phonon modes.
@@ VII.3.1 General Theoretical framework
Figure 9: Diagrammatic representation of the photon mediated interaction. (a) Mediated interaction to lowest order in $\Delta_{a}$ involving the exchange of a single photon. The solid orange lines represent the ground state propagator, the red lines depict the excited state propagator. As illustrated in the inset, the photon exchange process arises from a second-order process where an atom from the ground state is excited via a classical pump field (crossed dot with the black wiggly line), which then emits a photon (wavy blue line), which is absorbed in the reversed process from the inset. (c) The effective interaction between ground state atoms is denoted by $\mathcal{D}.$
@@
First, our main assumption is that the pump and cavity frequencies are far detuned from the excited state such that $\Delta_{a}$ is large. This is the so-called dispersive regime where the population of the excited state is small and one can write their Green’s function as $G_{e}(k)\simeq 1/\Delta_{a}$. The leading diagram in $1/\Delta_{a}$ for the photon-mediated interaction between the ground state atoms can then be drawn as in Fig. 9(a), which intuitively can be interpreted from left to right as: a ground state atom (orange line) is excited to the excited state (red line) by the classical pump field $\Omega$ (red cross with the black wiggly line). The excited state then decays to the ground state again by emitting a photon into cavity mode (wavy blue line). The cavity photon is subsequently absorbed by a second ground state atom, which is then excited. This atom finally decays back into the ground state atom via stimulated emission into the classical field. This completes the mediated interaction, as two ground state atoms effectively exchange a photon. As illustrated in Fig. 9(b), the processes highlighted by the box can be summarized into an effective interaction $\eta({\mathbf{r}})=\Omega({\mathbf{r}})\mathcal{G}({\mathbf{r}})/\Delta_{a}$ between a ground state atom and a photon. Thus, as for the other mediated interactions discussed in this perspective, the photon mediated interaction between the ground state atoms is quadratic in $\eta$. Finally, assuming that the dynamics of the cavity photon can be ignored when $\Delta_{c}$ is large so that its Green’s function is $G_{c}(k)\simeq 1/\Delta_{c}$ and taking into account the two processes in Fig. 9(c) gives
@@
for the photon mediated interaction between the ground state atoms. As described below, one reaches this expression for $\mathcal{D}(\mathbf{r},\mathbf{r}^{\prime})$, following a standard procedure for the Heisenberg equations of motion for the atom and photon field operators. Note that this interaction is infinite range reflecting that the cavity mode is assumed to cover the whole system. The photon-mediated interactions can be derived alternatively using a path-integral formalism as in Sec. IV.2. This approach has been employed for the study of supersolid phases of bosonic atoms inside two crossed optical cavities Lang et al. (2017); Piazza et al. (2013).
@@
$\left.+g_{eg}{{\hat{\psi}}_{e}^{\dagger}(\mathbf{r}){\hat{\psi}}_{e}(\mathbf{r})}\right]{\hat{\psi}}_{g}(\mathbf{r})+\hbar\left[\Omega(\mathbf{r})+\mathcal{G}(\mathbf{r})\hat{a}^{\dagger}\right]{\hat{\psi}}_{e}(\mathbf{r})$ (34)
@@
Here, we have made the identifications, ${\partial_{t}{\hat{\psi}}_{g}(\mathbf{r})=\partial_{t}{\hat{\psi}}}(\mathbf{r})$, $V({\mathbf{r}})=\Omega^{2}({\mathbf{r}})/\Delta_{a}$, $U({\mathbf{r}})=\mathcal{G}({\mathbf{r}})/\Delta_{a}$, $\hat{n}(\mathbf{r})={\hat{\psi}}^{\dagger}_{g}(\mathbf{r}){\hat{\psi}}_{g}(\mathbf{r})$, and $g_{gg}=g_{0}$. The above equations establish the coupled dynamics between matter and light.
@@
[…]
with $\mathcal{D}(\mathbf{r},\mathbf{r^{\prime}})$ given by Eq. 32.